OpenStudy (anonymous):

What is the solution to the equation 7^3x ≈ 9 ?

OpenStudy (anonymous):

7^3x≈9 ln7^3x≈ln9 3xln7≈ln9 x=ln9/3ln7 x=ln2/3

OpenStudy (anonymous):

how do i make this a decimal...? Im sorry i am just really bad at math can u show me?

OpenStudy (luffingsails):

If you are using a calculator, put in 2 divided by 3 then hit the equals key... then hit the natural log (or ln) button.

OpenStudy (luffingsails):

then entirely different answer <laugh>. enter 2, hit ln, then divide by 3.

OpenStudy (anonymous):

i have different choices... im confused on how to do this x = 0.376 x = 0.295 x = –0.295 x = –0.376

OpenStudy (anonymous):

Help please :(